# ExploraCoder: Advancing code generation for multiple unseen APIs via planning and chained exploration

### v1

@@ Abstract
Through training on publicly available source code libraries, large language models (LLMs) can invoke multiple encapsulated APIs to solve complex programming problems. However, existing models inherently cannot generalize to use APIs that are unseen in their training corpora. As libraries continuously evolve, it becomes impractical to exhaustively retrain LLMs with new API knowledge. This limitation hampers LLMs from solving problems which require newly introduced or privately maintained libraries. Human programmers often explore unfamiliar APIs by writing experimental code before invoking them for a more complex problem. Inspired by this behavior, we propose ExploraCoder, a training-free framework that empowers LLMs to invoke multiple unseen APIs in code solution by (1) planning a complex problem into several API invocation subtasks, and (2) exploring correct API usage through a novel chain-of-API-exploration. Concretely, ExploraCoder guides the LLM to iteratively generate several experimental API invocations for each simple subtask, where the promising execution experience are exploited by subsequent subtasks. This forms a chained exploration trace that ultimately guides LLM in generating the final solution. We evaluate ExploraCoder on Torchdata-Github benchmark as well as a newly constructed benchmark that involves more complex API interactions. Experimental results demonstrate that ExploraCoder significantly improves performance for models lacking prior API knowledge, achieving an absolute increase of 11.24% over naive RAG approaches and 14.07% over pretraining methods in pass@10. Moreover, the integration of a self-debug mechanism further boosts ExploraCoder’s performance on more challenging tasks. Comprehensive ablation and case studies provide further insights into the effectiveness of ExploraCoder. [^1] Code and data will be available at https://github.com/wykun-zju/ExploraCoder_paper.
@@ 1 Introduction
Library-oriented code generation refers to the automatic generation of code that leverages library APIs to solve specific programming problems (Zan et al., 2022; Liu et al., 2023). This task poses a challenge in recent AI research due to the complexity of diverse API usage and the interactions between different APIs (Alrubaye et al., 2019; Zan et al., 2024). Recent success of large language model (LLM), such as ChatGPT (OpenAI, 2022) and CodeLlaMA (Rozière et al., 2024), has demonstrated remarkable capability in various code generation tasks (Yan et al., 2024). Researchers have found that LLMs can effectively invoke APIs by pretraining on vast amounts of contemporary public libraries (Zan et al., 2023). However, a persistent challenge arises when the target API knowledge is sparse, outdated, or entirely unseen in the training data. This limitation hampers LLMs from problem solving that requires newly introduced or privately maintained libraries.
Prior work proposed to use continual pretraining (Gururangan et al., 2020) techniques to address this knowledge gap (Zan et al., 2022). But it is often impractical to exhaustively retrain LLMs since the code libraries are continuously evolving. And training data is especially hard to collect for newly introduced libraries. Another line of work adopts a naive retrieval-augmented generation (RAG) framework for unseen-library-oriented API invocations (Zan et al., 2023; Zhou et al., 2023; Liu et al., 2023), where the code generation process is divided into two phases: A retriever model retrieves relevant APIs from library documents. Then a generator model generates a code solution based on the problem description and retrieval results. However, they are mainly designed for simple API invocation tasks. Recent studies have shown that these existing approaches exhibit limited performance in complex programming tasks that require multiple API invocations (Zan et al., 2023; 2024; Ma et al., 2024b).
We propose to solve this issue by optimizing the RAG framework, as current solutions fall short on two fronts. On the one hand, it is hard for retrievers to map a comprehensive user requirement to a diverse range of APIs with various functionalities. On the other hand, existing generators lack the reasoning capability to handle intricate interactions among multiple APIs, especially when the models have not been trained on their usage knowledge.
Figure 1: An Overview of ExploraCoder Framework. ExploraCoder processes the given problem through Task Planning, API Recommendation, and Chain of API Exploration modules. The grey block in the bottom-left corner illustrates the detailed exploration process in the Chain of API Exploration. Finally, the processed results are used by a solution generator to generate final code solutions for the programming problem.
Exploratory programming (Sheil, 1986; Beth Kery & Myers, 2017) is a paradigm where programmers learn to do an unfamiliar task by actively experimenting with different possibilities using code. Specifically, when human programmers are tasked to solve a problem using an unfamiliar library. they would first learn the library’s capability from documents, then experiments with relevant API calls to gain hands-on usage experience.
@@
Inspired by this behavior, we propose ExploraCoder, a training-free framework aiming to facilitate LLM to invoke multiple unseen APIs through planning and a chain of API exploration (CoAE). As shown in Figure 1, given a complex programming problem, ExploraCoder first plans a series of simple API invocation subtasks based on library document. And a set of relevant APIs is recommended for each subtask. Then, a chained API exploration is carried out to iteratively experiments with various API invocations and pass on valuable usage experience to subsequent subtasks. This process forms an API exploration trace, which serves as a chain-of-thought (CoT) (Wei et al., 2022) instruction to guide the LLM in generating the final code solution.
We evaluate ExploraCoder on an existing multi-API benchmark (Torchdata-Github) and a newly constructed benchmark (Torchdata-Manual) that involves more complex API interactions. Experimental results demonstrate that ExploraCoder significantly improves performance for models lacking prior API knowledge, achieving an absolute increase of 11.24% over naive RAG approaches and 14.07% over pretraining methods in pass@10. Moreover, we find that the integration of a self-debug mechanism in CoAE further boosts ExploraCoder’s performance on more challenging tasks. Comprehensive ablation and case studies are conducted to provide further insights into the effectiveness of ExploraCoder.
@@
  • A novel iterative generation method. We propose a novel Chain-of-API-Exploration (CoAE) mechanism for step-wise code generation. In each intermediate step, CoAE generates partial API invocations based on limited API documentation and leverages execution signals to rectify coding decisions in real time.
  • An Exploratory-Programming-inspired framework. We design a unified framework, ExploraCoder, to integrate CoAE with other effective techniques. ExploraCoder is inspired by the exploratory programming paradigm observed in human developers(Beth Kery & Myers, 2017), who actively read API documents and gather more API usage experience from trial execution. Experimental results indicate the superiority of ExploraCoder’s design beyond the simple stacking of techniques.
  • A new multi-API-invocation benchmark. We manually construct a new benchmark, Torchdata-Manual, containing 50 complex, executable programming problems each involving 8-14 Torchdata API invocations. To the best of our knowledge, Torchdata-Manual features the longest API sequences among publicly reported execution-based library-oriented benchmarks.
  • Extensive experimental evaluation. We conduct extensive experiments in two unseen-library-oriented code generation benchmarks to validate the performance of ExploraCoder, demonstrating its superiority over SOTA multi-API code generation approaches.
@@ 2 Related Work
Code Generation with LLM. Code generation, the process of automatically producing program code from user specifications, has seen remarkable advancements with the advent of LLMs (Guo et al., 2024; Qwen, 2024; OpenAI., 2024). By training on vast code corpora, these models have demonstrated a profound ability to generate contextually accurate and executable code snippets from high-level descriptions.
Recent research has increasingly focused on enhancing LLM performance in handling complex coding tasks. A line of work attempts to adopt the chain-of-thought (Wei et al., 2022) approach from the NLP community, where LLMs are prompted to generate intermediate reasoning steps before arriving at the final answer. However, Self-Planning(Jiang et al., 2024) argues that the original idea of CoT does not reduce the difficulty of coding problems, and advocates for more effective planning mechanisms tailored specifically for code generation. Another line of work utilizes execution feedback to improve generated code(Madaan et al., 2023). For instance, Self-Repair(Olausson et al., 2024) proposes a three-step framework that involves generating initial code, reflecting on error messages, and subsequently repairing the code based on those reflections.
Library-Oriented Code Generation. Real-world programming problems often involve the use of external libraries, posing a challenge for LLM when tasked to invoke APIs that were not seen in the training data. An intuitive solution to this issue is to pretrain or continue pretrain on the new API data(Zan et al., 2022). However, the training process is often computationally expensive and time consuming, making this approach difficult to deploy in practice.
An alternative approach leverages RAG techniques. Most previous studies adopt a naive RAG framework, where a retriever directly retrieves relevant APIs and provides them as context for final solution generation. However, Zan et al. (2024) observed that such a framework struggles with complex problems requiring multiple API invocations. Some recent work puts efforts into improving existing RAG frameworks. For examples, Ma et al. (2024b) proposed a decomposed retrieval method, incorporating an inter-task LLM reranker to suggest k APIs for code generation. However, their approach primarily focuses on API retrieval, and its performance in generating multiple API invocations remains limited. Li et al. (2024) fine-tuned a task decomposer using proprietary enterprise data and integrated RAG with CoT techniques by putting programming subtasks in context as step-by-step instructions for the generator. However, their evaluation relies solely on lexical and human scoring metrics, leaving the execution performance of their method unexplored.
Unseen Library Benchmarks. Constructing unseen library benchmarks [^2] Also referred to as private library benchmark in some work is particularly challenging due to limited resources and the difficulty of verifying whether the knowledge was included in model training. Previous work has generally involved manually building small-scale benchmarks for mono-library invocation tasks. For example, Zan et al. (2023) lexically transformed API names in public library programming problems to build MonkeyEval and BeatNumEval. But the API functionalities are already exposed to LLMs. Zan also developed TorchdataEval with 50 simple API invocation problems. However, these benchmarks only address basic programming problems mostly involving 1-2 API invocations. Ma et al. (2024b) proposed a more advanced Torchdata benchmark, featuring 50 programming tasks adapted from GitHub client code, each involving 3–8 API invocations. But the benchmark remain relatively simple, as many tasks focus on basic file reading operations with 3–4 API calls, whereas real-world software development often involves far more complex API interactions (Kula et al., 2018; Bauer et al., 2012). This gap highlights the need for a more comprehensive and complex unseen library benchmark.
@@ 3.1 Task Definition
This work targets to addressing the task of library-oriented code generation (Zan et al., 2022), where the goal is to utilize library APIs to solve a programming problem. Formally, given a problem $\psi$ that specifies the user requirement and a set of API documents $\mathcal{A}$, a model $\theta$ generates solutions $p\sim\mathcal{P}_{\theta}(.|\psi,A)$.
Most code libraries provide essential information such as API signatures, descriptions, library overviews, and example usage code. In this paper, we assume the accessibility of this information from the provided documents. As shown in Figure 1, ExploraCoder will automatically identify relevant subset of APIs $\mathcal{\hat{A}}$ and accumulate useful invocation experience $\hat{\mathcal{E}}$, which are then used as augmenting signals to enhance the code generation process:
@@ 3.2 Planning for API invocation
Real-world programming problems often involve multiple composite operations (Yu et al., 2024), necessitating a plan for where and how APIs can contribute to problem-solving. Specifically, we need to outline several simple API invocation subtasks, upon which ExploraCoder will sequentially explore the correct API calls. Ideally, we aim to set the planning granularity to simple subtasks where each requires only 1–2 API invocations to be completed. Excessive granularity can make subtasks overly challenging, while incorrect segmentation—such as planning subtasks that cannot be addressed by any of the library APIs—can lead the model to generate numerous hallucinations along the line (Liu et al., 2024a; Tian et al., 2024).
However, the granularity and functional boundaries of library APIs are domain-specific, often falling out of distribution (OOD) of LLMs when the library is absent from the model’s training data. This misalignment poses a challenge in coordinating task planning with the typical usage patterns of these APIs.
To address this, we leverage the in-context learning capabilities of LLMs (Bareiß et al., 2022; An et al., 2023) by providing a condensed library overview and a small number of planner examples. This enables the LLMs to learn high-level usage patterns of the library. In this work, we prompt GPT-3.5-turbo-0125 to automatically summarize a piece of text $s$ from the library overview and extract few-shot planners $\mathcal{D}=\{\langle\psi_{j},\{t_{u}\}_{u=1}^{w_{j}}\rangle\}^{n_{\mathcal{D}}}_{j=1}$ from the provided code examples, where $\psi_{j}$ is the requirement description of the $j$-th code example, and $t_{u}$ is the explanation of $u$-th API invocation. Note that we do not leak any detailed API usage or benchmark-related knowledge to models (detailed in Appendix A.9). Now, we can plan $n$ API invocation subtasks for a given problem $\psi$:
@@ 3.3 API Recommendation
The API recommendation module serves to recommend relevant API documents $\mathcal{A}_{i}=\{a^{(1)},\dots,a^{(k)}\}$ for each API invocation subtask $t_{i}$ The API documents are transformed into a tabular style and then serve as the candidates for retrieval, where each row consists of the API import path, signature, and description. We first use a dense retriever to retrieve a initial subset of APIs by computing the similarity between $t_{i}$ and each $a_{j}$. More discussion of this module is provided in Appendix A.7.
@@
Then, we prompt LLM to re-rank and drop irrelevant ones from the APIs retrieval results for each subtask, providing the refined set $\{\mathcal{\tilde{A}}_{i}\}^{n}_{i=1}$ for Chain of API Exploration, where then the actually used APIs $\tilde{\mathcal{A}}_{\text{CoAE}}$ will be recorded. Meanwhile, we also conduct an inter-task reranking (Ma et al., 2024b) to recommend a subset $\tilde{\mathcal{A}}_{G}$ with the volume of $k_{G}$ APIs from a global perspective. In the final solution stage, we provide for the generator:
@@ 3.4 Chain of API Exploration
After preparing $n$ API invocation subtasks $\{\langle\tilde{\mathcal{A}}_{i},t_{i}\rangle\}_{i=1}^{n}$ through the previous two modules, the next step is to find the correct APIs and their appropriate usage for each subtask.
Previous work shows LLMs struggle to directly invoke multiple unseen APIs in a single run to solve a complex problem (Zan et al., 2024). The difficulty arises because LLMs tend to hallucinate the usage of unfamiliar APIs, and errors can propagate through APIs interactions, further compounding the issue. In contrast, when lacking knowledge of relevant APIs, programmers could apply an exploratory programming style, experimenting with code in a sandbox environment to accumulate experience about correct API usage. Inspired by this behaviour, we designed a Chain of API Exploration to help LLMs autonomously explore API usage and sequentially solve the tasks related to the problems. We now formalize the main steps in CoAE.
Experimental code generation. We encourage the LLM to generate $m$ diversified experimental code snippets for subtask $t_{i}$:
@@
where $s$ is the high-level library information from Section 3.2, and $\mathcal{E}_{1:i-1}$ is the accumulated invocation experience from subtasks prior to $t_{i}$. We define API invocation experience as the combination of a code snippet and its observed execution output, which will be elaborated in the next step. Each experimental code will attempt to solve the subtask by making different API invocations, and print out important information such as API’s returned object. Such feedback will be observed by LLM in the next step.
In our experiments, we use the example inputs provided by the programming problem as a hint, and prompt LLM to reuse or create its own test input for experiments. For a subtask that requires interactions with prior subtasks, the LLM adapts the code snippet from prior invocations to construct inputs for the currect API invocation.
Code execution and observation. A line of work execute candidate codes against prepared test cases to facilitate code ranking (Shi et al., 2022) or refinement (Zhang et al., 2023). However, quality testbeds are often inaccessible in production environments (Siddiq et al., 2023; Schäfer et al., 2023), particularly when developing new functionalities using newly introduced or unfamiliar libraries. Furthermore, creating test cases at runtime for the newly planned subtasks is even more challenging. Instead, we directly execute the experimental codes in a sandbox environment and capture its output. Specifically, given $t_{i}$ and $p_{i,j}$, the observation $o_{i,j}$ by the LLM consists of the codes’ executability $\delta$ , error message $\varepsilon$, and program output $\gamma$. We now can assemble $m$ candidate experience for $t_{i}$ as:
@@
Enhance experience exploitation by self-debugging. In our preliminary experiments, we found the experimental codes often fail to execute due to simple mistakes (e.g., missing import statements). Additionally, some challenging tasks require complex interactions with APIs from prior subtasks, which LLMs struggle to solve without execution feedback. preventing the observation of valuable insights into API behavior. And these failures could propagate along the exploration chain.
To address this, we prompt the LLM to autonomously repair failed codes when all candidate codes for a given task fail to execute, thereby enriching its execution experience. We report the effectiveness of ExploraCoder, both with and without self-debugging capabilities in Section 5.

### v2

@@ Abstract
Large language models face intrinsic limitations in coding with APIs that are unseen in their training corpora. As libraries continuously evolve, it becomes impractical to exhaustively retrain LLMs with new API knowledge. This limitation hampers LLMs from solving programming problems which require newly introduced or privately maintained libraries. Inspired by exploratory programming paradigm in human behavior, we propose ExploraCoder, a training-free framework that empowers LLMs to invoke multiple unseen APIs in code solution by (1) planning a complex problem into several API invocation subtasks, and (2) experimenting with correct API usage at intermediate steps through a novel chain-of-API-exploration. We conduct evaluation on program synthesizing tasks involving complex API interactions. Experimental results demonstrate that ExploraCoder significantly improves performance for models lacking prior API knowledge, achieving absolute increases of up to 11.99% over retrieval-based approaches and 17.28% over pretraining-based methods in pass@10.
ExploraCoder: Advancing Code Generation for Multiple Unseen APIs
via Planning and Chained Exploration
Yunkun Wang1thanks: Work is done during internship at Alibaba Group.,   Yue Zhang2,   Zhen Qin1,   Chen Zhi122footnotemark: 2, Binhua Li2,   Fei Huang2,   Yongbin Li2thanks: Corresponding authors.,   Shuiguang Deng 122footnotemark: 2 1Zhejiang University,   2Alibaba Group, {wangykun, zhenqin, zjuzhichen, dengsg}@zju.edu.cn {shiyu.zy, binhua.lbh, f.huang, shuide.lyb}@alibaba-inc.com
@@ 1 Introduction
Library-oriented code generation refers to the automatic generation of code that utilizes specified library’s APIs to solve programming problems (Zan et al., 2022; Liu et al., 2023). This task becomes particularly complex when the solution requires the integration of multiple APIs from the library, demanding not only knowledge of individual API functionalities but also an understanding of their interactions and dependencies (Alrubaye et al., 2019; Zan et al., 2024). Modern large language model (LLM), such as ChatGPT (OpenAI, 2022) and CodeLlaMA (Rozière et al., 2024), has demonstrated remarkable capability in generating API invocations using prior knowledge from pretraining stage (Zan et al., 2023). However, a significant challenge arises when the target API knowledge is sparse, outdated, or entirely unseen in the training data. This limitation hampers LLMs from problem solving that requires newly introduced or privately maintained libraries.
Prior work proposed to use continual pretraining (Gururangan et al., 2020) to address this knowledge gap (Zan et al., 2022). But this is often impractical due to the scarcity of training data for new libraries and the substantial costs of retraining LLMs. Another line of work adopts a standard retrieval-augmented generation (RAG) framework for unseen API invocations (Zhou et al., 2023; Zan et al., 2023; Liu et al., 2023), where LLM acquires API knowledge from retrieving the library documentation. While effective for simple API invocation tasks, these methods struggle with complex scenarios requiring multiple API invocations (Zan et al., 2023, 2024; Ma et al., 2024).
More recent studies propose to address complex API invocation tasks by improving document-retrieval (Ma et al., 2024) and preactively planning coding steps (Li et al., 2024). However, they overlook the challenge posed by the potential ambiguities in the API documentation. Some work adopts iterative or agentic workflow (Olausson et al., 2024; Yao et al., 2022; Zhu et al., 2024) to reactively plan for retrieval and debugging, however, the end-to-end code construction could still expose the limitations of LLMs in coordinating multi-API interactions.
@@
When coding with an unfamiliar library, experienced developers would adpot an Exploratory Programming paradigm (Sheil, 1986; Beth Kery and Myers, 2017). This involves first understanding the library’s capabilities through documentation to devise a broad plan, and then actively experimenting with individual API calls to gain practical experience, ultimately leading to a correct code solution. Inspired by this behavior, we propose ExploraCoder, a training-free framework aiming to facilitate LLM to invoke multiple unseen APIs. As shown in Figure 1, given a complex programming problem, ExploraCoder begins by planning a series of simpler API invocation subtasks based on library documentation. For each subtask, it recommends a set of candidate APIs. Subsequently, a Chain of API Exploration (CoAE) is performed, iteratively experimenting with various subtask-wise API invocations while passing valuable usage insights to the plannings of subsequent subtasks. This process forms an API exploration trace, which facilitates the LLM in deriving the final solution.
Evaluating unseen-library-oriented code generation requires an unexposed library. Prior works (Zan et al., 2023; Ma et al., 2024) have created simple benchmarks using the Torchdata library, as it strikes a balance between minimum client code use for researcher’s problem designing and limited exposure to modern LLMs. However, benchmarking multiple unseen-API tasks remains underexplored. To address this gap and better reflect real-world complex programming challenges, we constructed a new Torchdata-based benchmark, Torchdata-Manual, featuring complex multi-API problems. Experimental results on our Torchdata-Manual and an existing Torchdata-Github benchmarks demonstrate that ExploraCoder significantly improves performance for models lacking prior API knowledge, achieving absolute gains of up to 11.99% over various retrieval-based approaches and 17.28% over pretraining methods in pass@10. Moreover, we find the integration of an intermediate self-debug mechanism further boosts ExploraCoder’s performance on more challenging tasks.
@@
  • We propose ExploraCoder, a unified framework that incorporates unseen API knowledge from documentation into a novel step-wise code generation method, Chain-of-API-Exploration. By leveraging this framework, LLMs can plan based on library documentation and actively experiment with APIs in intermediate steps, mirroring the Exploratory Programming paradigm employed by human developers.
  • We construct Torchdata-Manual, a new library-oriented benchmark that, to the best of our knowledge, features the highest number of API invocations per task among publicly reported executable library-oriented benchmarks. The code and data are available at  https://github.com/greenlight2000/ExploraCoder.
  • Experimental results and case studies on ours and an existing benchmark demonstrate ExploraCoder’s superior performance on multi-API tasks compared to competitive baselines.
Figure 1: An Overview of ExploraCoder Framework. ExploraCoder processes the given problem through Task Planning, API Recommendation, and Chain of API Exploration modules. The gray block in the bottom-left corner illustrates the detailed inner process in the Chain of API Exploration. Finally, the processed results are used by a solution generator to generate final code solutions for the programming problem.
@@ 2 Related Work
Complex Code Generation with LLM.
Code generation, the process of producing code from NL specifications, has seen remarkable advancements with LLMs (OpenAI., 2024; Yan et al., 2024). Recent research has increasingly focused on LLMs tackling complex coding tasks, such as competition (Li et al., 2022), library-oriented (Bogomolov et al., 2024), and repo-level (Jimenez et al., 2024) problems. One prominent paradigm leverages chain-of-thought (Wei et al., 2022) to plan for intermediate steps before complex code generation, whereas its effectiveness diminishes when high quality plans cannot be derived (Jiang et al., 2024). Another direction proposes to debug after the generation of code (Olausson et al., 2024), but they typically require accessibility of test cases. ExploraCoder distinguishes itself by applying an intermediate code construction that leverages executability signals to rectify coding plans at steps in real time.
Library-Oriented Code Generation.
Real-world programming problems often involve the use of external libraries, posing a challenge for LLMs to invoke APIs unseen from training data. Continued pretraining on the new API data, though intuitive, is often impractical due to its complexity and cost. Most prior studies adopt a naive RAG framework as an alternative to incorporate APIs knowledge from library documentation (Zhou et al., 2023; Zan et al., 2023). But they struggle with more complex problems that require multiple API invocations (Zan et al., 2024). Recent studies have attempted to improve the RAG frameworks. For examples, CAPIR (Ma et al., 2024) proposed a decomposed the retrieval process to identify accurate API docs. EpiGen (Li et al., 2024) makes preactive NL plans for one-pass code generation. These works mainly focus on preprocessing relevant API context, while overlooking the reasoning limitation of LLMs in multi-API interactions, and the challenge posed by the potential ambiguity in API documentation.
Unseen Library Benchmarks.
Constructing unseen library benchmarks is particularly challenging, as libraries new enough to have limited exposure to modern LLMs often lack the rich client code needed for developing complex problems. Previous work has generally turned to a Torchdata library to manually build small-scale API invocation benchmarks.
Zan et al. (2023) constructed TorchdataEval mostly involving 1-2 simple API invocations. Ma et al. (2024) introduced 50 multi-API programming tasks adapted from Torchdata client code from Github, each involving 3–8 API invocations. However, some of its tasks remain relatively simple and real-world development often involves more API interactions (Kula et al., 2018; Bauer et al., 2012). This gap highlights the need for a more complex unseen library benchmark.
@@ 3.1 Task Definition
This work addresses the task of library-oriented code generation (Zan et al., 2022). Formally, given a problem $\psi$ that specifies the user requirement and a library API documentation $\mathcal{A}$, a model $\theta$ generates code solutions $p\sim\mathcal{P}_{\theta}(.|\psi,A)$.
Most code libraries provide basic information about their APIs, such as API signatures, descriptions, and high-level library overviews with minimum example usage code. In this paper, we assume the accessibility of this information from API documents. As shown in Figure 1, ExploraCoder will automatically identify relevant subset of APIs $\mathcal{\hat{A}}$ and accumulate useful experience of intermediate API invocations $\hat{\mathcal{E}}$, both of which are then used as augmenting signals to generate the final solution:
@@ 3.2 Planning for API invocation
Real-world programming problems often involve composite operations (Yu et al., 2024), necessitating a plan for where and how APIs can contribute to problem-solving. Specifically, we need to outline several API-related subtasks, upon which ExploraCoder will sequentially explore the correct API calls. Ideally, we aim to set the planning granularity to simple subtasks where each requires only 1–2 unseen API invocations. However, the functional granularity of APIs is domain-specific, often falling out of distribution (OOD) of LLMs when the library is absent from their training data, posing a challenge in aligning task planning with typical API usage patterns.
To address this, we leverage the in-context learning capabilities of LLMs (An et al., 2023) by providing a condensed library overview and a small number of planner examples. This enables the LLMs to learn high-level usage patterns of the library without needing to know all its APIs. In this work, we prompt GPT-3.5-turbo-0125 to automatically summarize a piece of text $s$ from the library overview and extract few-shot planners $\mathcal{D}=\{\langle\psi_{j},\{t_{u}\}_{u=1}^{w_{j}}\rangle\}^{n_{\mathcal{D}}}_{j=1}$ from the provided code examples, where $\psi_{j}$ is the requirement of the $j$-th code example, and $t_{u}$ is the explanation of $u$-th API invocation. Note that we do not leak any detailed API usage or benchmark-related knowledge to models (detailed in Appendix A.8). Now, we can plan $n$ API-related subtasks for a given problem $\psi$:
@@ 3.3 API Recommendation
The API recommendation module serves to recommend relevant API documents $\mathcal{A}_{i}=\{a^{(1)},\dots,a^{(k)}\}$ for each API invocation subtask $t_{i}$ We process the documents into tabular retrieval pool, where each row consists of the API import path, signature, and description. We first use a dense retriever to retrieve an initial set of APIs by computing the similarity between $t_{i}$ and each $a_{j}$.
@@
Then, we prompt LLM to re-rank and drop irrelevant APIs for each subtask, providing a refined subset $\{\mathcal{\tilde{A}}_{i}\}^{n}_{i=1}$ for Chain of API Exploration, where then the actually used APIs $\tilde{\mathcal{A}}_{\text{CoAE}}$ will be recorded. Meanwhile, we also conduct an inter-task reranking (Ma et al., 2024) to recommend $k_{G}$ APIs $\tilde{\mathcal{A}}_{G}$ from a global perspective. In the final solution stage, we provide for the generator:
@@ 3.4 Chain of API Exploration
Previous work shows LLMs struggle to directly invoke multiple unseen APIs in a single run (Zan et al., 2024). The challenge arises from LLMs’ tendency to hallucinate unfamiliar APIs usage. Hallucinations in early decoding step could compromise subsequent API calls due to the autoregressive nature of LLM, further compounding the error.
In contrast, when lacking knowledge of relevant APIs, developers could adopt an exploratory programming paradigm, actively experimenting with partial code in a sandbox environment to accumulate correct API usage experience. Inspired by this behavior, we designed a Chain of API Exploration (CoAE) to sequentially explore API usage and solve the $n$ subtasks $\{\langle\tilde{\mathcal{A}}_{i},t_{i}\rangle\}_{i=1}^{n}$. We now formalize the main steps in CoAE.
Experimental code generation.
We prompt the LLM to generate $m$ diversified experimental code snippets for intermediate subtask $t_{i}$:
@@
where $s$ is the high-level library information from Section 3.2, and $\mathcal{E}_{1:i-1}$ is the accumulated invocation experience from prior subtasks that could further enhance the preactive planning of $t_{i}$. We define API invocation experience as the combination of an intermediate code snippet and its execution output, which is elaborated in the next paragraph. Each experimental code will attempt to solve the subtask by making different API invocations, and print out valuable usage knowledge. Such feedback will be observed by LLM in the next step.
Code execution and observation.
At each subtask, LLM is encouraged to print out insightful information to expand API usage knowledge, such as format of the current API returned object that could be used as input in other subtasks. We capture the output from directly executing the experimental code in a sandbox environment. Specifically, given $t_{i}$ and $p_{i,j}$, the observation $o_{i,j}$ by the LLM consists of the codes’ executability $\delta$ , error message $\varepsilon$, and program output $\gamma$. We now can assemble $m$ candidate API invocation experience for $t_{i}$ as:
@@
Experience exploitation by intermediate self-debugging.
In our preliminary experiments, we found experimental codes often fail to execute due to simple mistakes (e.g., missing import statements). Additionally, some challenging subtasks require complex API interactions with prior subtasks, which LLMs struggle to solve. This hinders the acquisition of additional API usage insights, and the intermediate failures could potentially degenerate the performance of exploration chain. To address this, we prompt the LLM to debug the codes when all candidate codes for a given subtask fail to execute, thereby enhancing its usage experience. We report the effectiveness of ExploraCoder, both with and without debugging mechanism.
Experience selection strategy.